# StrandHead: Text to Strand-Disentangled 3D Head Avatars Using Hair Geometric Priors

### v1

@@ Abstract
While haircut indicates distinct personality, existing avatar generation methods fail to model practical hair due to the general or entangled representation. We propose StrandHead, a novel text to 3D head avatar generation method capable of generating disentangled 3D hair with strand representation. Without using 3D data for supervision, we demonstrate that realistic hair strands can be generated from prompts by distilling 2D generative diffusion models. To this end, we propose a series of reliable priors on shape initialization, geometric primitives, and statistical haircut features, leading to a stable optimization and text-aligned performance. Extensive experiments show that StrandHead achieves the state-of-the-art reality and diversity of generated 3D head and hair. The generated 3D hair can also be easily implemented in the Unreal Engine for physical simulation and other applications.
Figure 1: We propose StrandHead, a text-driven framework for generating strand-disentangled 3D head avatars that feature high-fidelity facial details and strand-based hair. By accurately capturing the internal geometry of hair strands, our approach seamlessly supports flexible hairstyle transfer and editing, as well as physics-based rendering and simulation.
@@ 1 Introduction
Creating 3D head avatars is crucial for many applications including digital telepresence, gaming, movies, and AR/VR. Traditional approaches heavily depend on manual efforts which is time-consuming and labor-intensive. Recent advances [55, 84, 93, 79] achieve automatic 3D head generation with the learning-based paradigm. Although these methods provide reliable 3D head geometry, they require high-cost 3D training data, which significantly constrains their generalization capabilities and limits the potential applications.
@@
With the rapid development of text-driven generation methods [57, 50], creating 3D head avatars from given prompts becomes possible. Recent studies [1, 12, 99, 89] build upon the standard text-to-3D generation frameworks, achieving low-cost and high-fidelity performance by integrating with 3D parametric head models [4, 26, 82]. While focusing on facial geometry and texture modeling, these methods utilize holistic meshes or NeRF [39] to represent 3D haircut, failing to capture the internal geometric structure of hair strands, i.e., 3D curves. This limitation not only significantly reduces the realism of the generated avatars but also makes them incompatible with existing strand-based rendering and simulation systems [10, 2, 9].
@@
To model 3D haircuts, recent efforts achieve strand-level hairstyle reconstruction [58, 66, 94] or generation [97, 13] using VAEs or parametric models. Despite their impressive performance, they require constrained multi-view images or manually latent space searching, failing to freely create 3D hair under semantics of textual prompts. The most recent work HAAR [67] pioneers text-to-strand generation by training a text-conditioned hair map diffusion model. However, due to the high cost and limited scale of available 3D hair data, HAAR struggles to generate hairstyles outside of the training set. In addition, HAAR does not model hair texture or geometry that adapts to a specific generated human head, which limits its applicability in downstream tasks.
In this paper, we propose StrandHead, a novel framework that generates strand-disentangled 3D head avatars from text descriptions. As illustrated in Fig. 1, the generated head avatars exhibit detailed 3D faces with high-fidelity hair strands and diverse 3D haircuts, enabling seamless and flexible haircut transfer, editing, and physical-based simulation. Without requiring any hair training data, we create 3D hair by distilling the 2D generative diffusion models. This is achieved by employing a series of key insights on hair geometric priors. First, inspired by the cylindrical structure of hair strands [19, 35], we propose a novel differentiable prismatization algorithm. This algorithm facilitates the efficient, flexible, and differentiable transformation of strands into watertight prismatic meshes with customizable thickness and defined lateral edges, enabling the use of mesh-based renderers, loss functions, and models for detailed strand-level modeling tasks. Second, leveraging the distribution patterns of 3D hair orientation and curvature, we propose reliable consistency regularization losses for robust generation. These losses enhance the cosine similarity between neighboring strand orientations and constrain the global hair curvature. Extensive experiments demonstrate that StrandHead outperforms the state-of-the-art (SOTA) methods in both head and hair generation tasks, and supports flexible haircut transfer and editing, as well as physical-based rendering and simulation.
@@
  • We propose StrandHead, a novel framework for generating high-fidelity 3D head avatars with disentangled strand-based hair. To the best of our knowledge, StrandHead is the first work to generate strand-disentangled 3D head from pre-trained 2D diffusion models.
  • We propose a differentiable prismatization algorithm that converts hair strands into watertight prismatic meshes, facilitating the use of mesh-based renderers, losses, algorithms, and models in strand-level deep learning tasks.
  • Inspired by statistical hair features, we propose orientation consistency and curvature regularization losses to supervise both local and global hair shape, enabling reliable and realistic hairstyle generation.
@@ 2 Related Work
Text-to-3D General Object Generation. Inspired by the success of text-to-image (T2I) generation [43, 53, 54, 59, 57, 91], many studies have explored using pre-trained vision-language models [52, 57] to achieve text-guided zero-shot 3D content generation. Early methods [40, 38, 61, 18, 71] employ the CLIP model [52] to supervise the alignment of 3D representations with prompts. DreamFusion [50] introduces the Score Distillation Sampling (SDS) loss, significantly enhancing the fidelity of generated 3D content by leveraging more powerful pre-trained diffusion models [57]. Subsequent works further advance text-driven 3D generation by improving 3D representations [5, 6, 85], optimization strategies [29, 70, 86], SDS loss [75, 28], and diffusion models [27, 51, 65]. Despite these advancements, current approaches focused on general content generation do not fully leverage the extensive prior knowledge of human heads and hair, limiting their ability to generate realistic, high-quality 3D head avatars with strand-based hair.
@@
Text-to-3D Head Avatar Generation. CLIPFace [1], T2P [92] and Describe3D [78] pioneer zero-shot text-driven 3D head generation by combining CLIP [52] with 3D parametric head models [4, 26, 8, 82]. With the introduction of SDS loss [50], DreamFace [90] and FaceG2E [80] leverage pre-trained T2I diffusion models to greatly improve generation fidelity within specific domains. HeadEvolver [72] enhances the expressiveness of head mesh deformations by introducing vector fields, while HumanNorm [17], HeadSculpt [12], and HeadArtist [30] employ DMTet [63] instead of traditional meshes to capture geometric details. HeadStudio [99] incorporates 3DGS [20] to produce realistic and animatable 3D heads. Despite these advancements, these approaches treat the head and hair as a holistic model, limiting support for downstream applications such as hairstyle transfer and editing. To enable flexible head-hair-disentangled generation, TECA [89] independently represents the head with mesh and the hair with NeRF [39]. However, these methods focus only on the realistic external appearance of hair without modeling its internal geometric structure, restricting their application for strand-level physics-based rendering and simulation.
Strand-Based Hair Creation. Automatic strand-based hair modeling has attracted substantial attention from both industry [10, 2, 9] and academia [87, 34, 3, 14], as it enables downstream physics-based rendering and simulation Such methods can be roughly divided into two categories: (1) reconstructing hair from input images or videos and (2) generating hair based on control conditions. Works on the former one [15, 96, 60, 42, 21, 58, 76, 83, 64, 66, 94, 33, 88, 77, 69, 95, 98, 36] relies on classical 3D reconstruction frameworks [62, 46, 37, 39, 20] and incorporates hair-specific features [45] to realize accurate 3D strand reconstruction. However, these methods depend on constrained multi-view images, making hair creation costly and challenging in less controlled settings. Recently, 3D hair generation approaches [97, 13] have emerged using GANs or parametric model, but they struggle with controllable generation. The most relevant work to ours is HAAR [67], which pioneers text-guided strand generation using diffusion models. However, HAAR relies on high-cost 3D hair data which is limited in diversity, restricting its ability to generate novel haircut beyond the training data.
Compared to previous methods, StrandHead generate 3D heads with fine geometry and lifelike textures, as well as realistic textured strand-based hairstyles, and the entire framework does not require any hair training data. We summarize the main differences between our approach and related approaches in Tab. 1.
Task Method Head-Hair- Decoupled Strand-Based Hair Geometry & Texture No Training Data
Text-to-Head [12, 30, 99, 17] [89]
[…]
@@
Figure 2: Strandhead consists of two stages: (a) Under the constraints of the human-specific diffusion model and the FLAME-volving prior loss, StrandHead generates a detailed and reasonable bald head. (b) By introducing a differentiable prismatization algorithm, orientation consistency loss and curvature regularization loss inspired by hair geometric priors, StrandHead achieves diverse and realistic strand-accurate hair creation without any requiring hair training data.
@@ 3.1 Preliminaries
FLAME [26] is a 3D parametric head model that represents the human head shape, pose, and expression using a compact set of parameters. Given a set of shape parameters $\beta$, pose parameters ${\theta}_{\text{head}}$, and expression parameters $\psi_{\text{exp}}$, FLAME can create a 3D head mesh ${M}$.
Score Distillation Sampling (SDS) [50] uses a pre-trained T2I model to guide the optimization of a 3D representation based on the input text. For an input text prompt $y$, a 3D representation parameterized by $\theta$, and a T2I diffusion model with parameters $\phi$, the SDS loss is defined as follows:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\phi}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],\vspace{-0.3cm}$ (1)
where $x=g(\theta)$ is the image rendered from $\theta$ by a differentiable renderer $g$, $t$ is the time step, $x_{t}=x+\epsilon$ is a noised version of $x$, ${\epsilon}_{\phi}(x_{t};y,t)$ is the denoised image, and $w(t)$ is a weighting function. For simplicity, we omit $w(t)$ in the following formulas.
DMTet [63] is a hybrid representation that marrys a signed distance function (SDF) and a differentiable Marching Tetrahedral layer to convert implicit SDF to an explicit mesh. Thanks to multi-resolution hash encoding [41], the SDF can be implemented in a lightweight manner.
Neural Scalp Textures (NST) [58] is an efficient hair representation where each pixel stores a feature vector conveying the shape information of a single strand at the corresponding scalp location. With the help of a pre-trained hair strand generator [58] $G$, the low-dimensional 2D neural scalp texture $T$ can be decoded into high-dimensional, dense 3D strand polylines $S=\{s^{i}\}_{i=1}^{N_{s}}$, where $N_{s}$ is the number of strand polylines, and $s^{i}=\{p^{i}_{j}\}_{j=1}^{N_{p}}$ is a polyline formed by $N_{p}$ 3D points. This process can be formulated as follows: $S=G(T)$.
@@ 3.2 Overview
Given a text prompt, StrandHead aims to create a high-quality 3D head avatar with strand-disentangled textured hair. Fig. 2 provides an overview of our pipeline. We first produce a high-fidelity and reasonable 3D bald head (Sec. 3.3). Then, we generate realistic, reasonable strand-based hair with a lifelike appearance by introducing a differentiable prismatization algorithm, two prior-driven losses and strand-aware texture field (Sec. 3.4).
@@ 3.3 Bald Head Generation
Head Geometry Modeling. To sculpt detailed geometry, we follow HeadArtist [30] and HeadSculpt [12] using DMTet [63] as our 3D representation and initialize it with the FLAME mesh $M_{\text{init}}$. After initialization, we leverage Nvdiffrastt [22] and SDS loss to model the geometry and texture of the human head in a decoupled manner. We employ a series of human-specialized diffusion models proposed by HumanNorm [17] to further improve the fidelity of the generated head. Specifically, we optimize the head DMTet parameterized by ${\theta}_{\text{h}}$ using the following losses:
$\nabla_{{\theta}_{\text{h}}}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{h}}_{t};y_{\text{h}},t)-\epsilon)\frac{\partial d^{\text{h}}}{\partial{\theta}_{\text{h}}}\right],\vspace{-0.5cm}$ (3)
Additionally, inspired by Barbie [68], we impose a FLAME-evolving prior loss to prevent unnatural geometry (e.g., overly large ears or overly pointed head tops). In particular, we freeze the FLAME shape parameters $\beta$ but improve $M_{\text{init}}$ to evolvable $\hat{M}_{\text{init}}$ by adding learnable vertex-wise offsets, and periodically fit $\hat{M}_{\text{init}}$ to the head DMTet every $\delta$ iterations, enabling the learned offsets to model complex facial features (e.g., beards and wrinkles, see the Supp. Mat. for loss details). During evolution, $\hat{M}_{\text{init}}$ gradually captures rich geometric details, and provide a reliable and diverse head prior for geometry generation using the following formula:
$\mathcal{L}_{\text{prior}}=\sum_{p\in P}\left\|s_{\theta_{\text{h}}}(p)-s_{\hat{M}_{\text{init}}}(p)\right\|_{2}^{2},\vspace{-0.4cm}$ (4)
where $s_{\theta_{\text{h}}}$ and $s_{\hat{M}_{\text{init}}}$ are the SDF functions of the head DMTet and $\hat{M}_{\text{init}}$, and $P$ is a set of randomly sampled points. In summary, the loss function for optimizing the head geometry is as follows:
$\mathcal{L}_{\text{head-geo}}=\mathcal{L}^{\text{hn}}_{SDS}+\mathcal{L}^{\text{hd}}_{SDS}+\lambda_{\text{prior}}\mathcal{L}_{\text{prior}}.\vspace{-0.2cm}$ (5)
Head Texture Modeling. Given the generated head geometry generated, we fix it and utilize a texture field ${\psi}_{\text{h}}$ which maps a query position to its color to generate head appearance. Specifically, we construct this field using a MLP network with multi-resolution hash encoding [41], and we optimize it with the following loss function:
$\nabla_{{\psi}_{\text{h}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h}}_{t};n^{\text{h}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h}}}{\partial{\psi}_{\text{h}}}\right],\vspace{-0.2cm}$ (6)
$\nabla_{{\psi}_{\text{h}}}\mathcal{L}^{\text{hc}}_{MSDS}=\mathbb{E}_{t,\epsilon}\left[(H(c^{\text{h}}_{t};n^{\text{h}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h}}}{\partial{\psi}_{\text{h}}}\right],\vspace{-0.2cm}$ (7)
where $H(\cdot)$ denotes the multi-step operation function.
3.4 Strand-Based Hair Generation
Hair Initialization. To achieve accurate and diverse hair initialization, we utilize ChatGPT [44] to select the 20 most representative hairstyles from the USC-HairSalon Dataset [15]. We then use the following loss to optimize NST to fit the selected hairstyle:

### v3

# StrandHead: Text to Hair-Disentangled 3D Head Avatars Using Human-Centric Priors

@@ Abstract
While haircut indicates distinct personality, existing avatar generation methods fail to model practical hair due to the data limitation or entangled representation. We propose StrandHead, a novel text-driven method capable of generating 3D hair strands and disentangled head avatars with strand-level attributes. Instead of using large-scale hair-text paired data for supervision, we demonstrate that realistic hair strands can be generated from prompts by distilling 2D generative models pre-trained on human mesh data. To this end, we propose a meshing approach guided by strand geometry to guarantee the gradient flow from the distillation objective to the neural strand representation. The optimization is then regularized by statistically significant haircut features, leading to stable updating of strands against unreasonable drifting. These employed 2D/3D human-centric priors contribute to text-aligned and realistic 3D strand generation. Extensive experiments show that StrandHead achieves the state-of-the-art performance on text to strand generation and disentangled 3D head avatar modeling. The generated 3D hair can be applied on avatars for strand-level editing, as well as implemented in the graphics engine for physical simulation or other applications. Project page: https://xiaokunsun.github.io/StrandHead.github.io/.
Figure 1: Given (a) input prompts, StrandHead generates (b) realistic 3D head avatars featuring strand-level attributes and (c) 3D hair strands by utilizing human-specific 2D generative priors and 3D hair strand geometric priors. By precisely capturing the internal geometry of hair strands, our approach enables seamless and flexible (d) hairstyle transfer and (e) editing, as well as (f) physics-based simulation.
@@ 1 Introduction
Creating 3D head avatars is crucial for many applications including digital telepresence, gaming, movies, and AR/VR. Hairstyles reflecting personal characteristics greatly impact the fidelity and realism of digital humans. Traditional methods rely heavily on manual effort, making them time-consuming and labor-intensive. Recent advances [52, 79, 88, 75] achieve automatic 3D head generation with the research paradigm based on supervised learning. However, these methods require costly 3D training data, which significantly limits their generalization capabilities and restricts potential applications.
@@
With the rapid development of text-driven generation methods [54, 47], creating 3D head avatars from given prompts without requiring 3D-text datasets becomes possible. Recent studies [1, 12, 94, 84] unlock the potential of 2D diffusion models in modeling 3D head avatars by integrating them with 3D head priors [4, 26, 77]. HumanNorm [17] further improves domain-specific fidelity by fine-tuning diffusion models on high-quality 3D human meshes. While focusing on facial geometry and texture modeling, these methods utilize holistic meshes or NeRF [38] to represent 3D haircuts, failing to capture the internal geometric structure of hair strands, i.e., 3D curves. This limitation not only significantly reduces the realism of the generated avatars but also makes them incompatible with strand-based applications and simulation systems [10, 2, 9].
@@
To accurately model 3D haircuts, recent efforts achieve strand-level hairstyle reconstruction [55, 64, 89] or generation [92, 13] using VAEs or parametric models. Despite their impressive performance, these methods require constrained multi-view images or manual latent space searching, preventing them from freely creating 3D hair based on user-friendly text. The most recent work, HAAR [65] pioneers text-to-strand by training a text-conditioned hair map diffusion model. However, due to its reliance on large-scale and costly paired data including 9825 haircuts and descriptions, HAAR faces challenges in generating diverse hairstyles outside the training set. Additionally, HAAR overlooks hair texture and geometry that adapts to specific head shapes, further limiting its practical applications.
So can we use powerful human-specific 2D generative priors to create realistic and diverse strand-based hair from text? We address this meaningful and challenging problem by developing StrandHead, a novel framework that generates high-fidelity 3D head avatars with strand-accurate hair from prompts. As illustrated in Fig. 1, the generated head avatars feature diverse and realistic hair strands, enabling seamless and flexible strand-level transfer, editing, and physics-based simulation. Instead of using large-scale hair-prompt paired data, we achieve text-aligned 3D hair strand generation by effectively utilizing human-specific 2D generative priors and 3D hair geometric priors.
Specifically, we propose a novel differentiable prismatization algorithm inspired by the cylindrical structure of hair strands. This algorithm can efficiently and differentiably convert strands into watertight prismatic meshes, which enables smooth backpropagation of gradients from 2D diffusion models pre-trained on human mesh data to 3D hair strands using mesh-based differentiable renderers, thereby making it possible to use 2D generative priors for modeling 3D hairstyles. Then we propose two effective losses inspired by the distribution patterns of 3D hair geometric features to further regularize the hair shape. These rich 2D/3D human-centric priors work together to achieve text-aligned and realistic 3D hair strand generation. Extensive experiments demonstrate that StrandHead outperforms the state-of-the-art (SOTA) methods in both head and hair generation tasks, and supports flexible haircut transfer and editing, as well as physical-based rendering and simulation.
@@
  • We propose StrandHead, a novel framework for generating realistic 3D head avatars with strand-level attributes. To the best of our knowledge, StrandHead is the first work to generate 3D hair strands by distilling human-specific 2D diffusion models.
  • We propose a differentiable prismatization algorithm that converts hair strands into watertight prismatic meshes. This ensures stable gradient flow from 2D generative priors to 3D hair strand representation, thereby achieving reliable end-to-end strand-based hair optimization.
  • Inspired by statistical 3D hair geometric features, we introduce two simple but solid regularization losses to supervise both local and global hair shapes, enabling reasonable and realistic hairstyle generation.
@@ 2 Related Work
Text-to-3D General Object Generation. Inspired by the success of text-to-image (T2I) generation [42, 50, 51, 56, 54, 86], many studies have explored using pre-trained vision-language models [49, 54] to achieve text-guided 3D content generation without large-scale 3D-text paired data. Early methods [39, 37, 58, 18, 69] employ the CLIP model [49] to supervise the alignment of 3D representations with prompts. DreamFusion [47] introduces the Score Distillation Sampling (SDS) loss, significantly enhancing the fidelity of generated 3D content by leveraging more powerful pre-trained diffusion models [54]. Subsequent works further advance text-driven 3D generation by improving 3D representations [5, 6, 80], optimization strategies [29, 68, 81], SDS loss [71, 28], and diffusion models [27, 48, 62]. Despite these advancements, current approaches focused on general content generation do not fully leverage the extensive prior knowledge of human heads and hair, limiting their ability to generate realistic, high-quality 3D head avatars with strand-based hair.
@@
Text-to-3D Head Avatar Generation. CLIPFace [1], T2P [87] and Describe3D [74] pioneer zero-shot text-driven 3D head generation by combining CLIP [49] with 3D parametric head models [4, 26, 8, 77]. With the introduction of SDS loss [47], DreamFace [85] and FaceG2E [76] leverage pre-trained T2I diffusion models to greatly improve generation quality within specific domains. HeadEvolver [70] enhances the expressiveness of head mesh deformations by introducing vector fields, while HeadSculpt [12], and HeadArtist [30] employ DMTet [60] instead of traditional meshes to capture geometric details. HeadStudio [94] incorporates 3DGS [20] to produce realistic and animatable 3D heads. HumanNorm [17], further enhances human-specific fidelity by fine-tuning 2D diffusion models on high-quality human mesh data. Despite these advancements, these approaches treat the head and hair as a holistic model, limiting support for downstream applications such as hairstyle transfer and editing. To enable flexible head-hair-disentangled generation, TECA [84] independently represents the head with mesh and the hair with NeRF [38]. However, these methods focus only on the realistic external appearance of hair without modeling its internal geometric structure, restricting their application for strand-level editing and physics-based simulation.
Task Method
Head-Hair-
Decoupled
Strand-Based
Hair
Geometry &
Texture
No Large-Scale
3D-Text Paired Data
Text-to-Head
[12, 30, 94, 17]
[84]
Text-to-Hair
[62, 80, 28, 48]
[65]
Text-to-Head-Hair
Ours
@@
Strand-Based Hair Creation. Automatic strand-based hair modeling has garnered significant attention from both industry [10, 2, 9] and academia [82, 33, 3, 14], as it enables downstream physics-based rendering and simulation Such methods can be broadly categorized into two types: (1) reconstructing hair from input images or videos and (2) generating hair based on control conditions. Works on the former one [15, 91, 57, 41, 21, 55, 72, 78, 61, 64, 89, 32, 83, 73, 67, 90, 93, 35] relies on classical 3D reconstruction frameworks [59, 45, 36, 38, 20] and incorporates hair-specific features [44] to realize accurate 3D strand reconstruction. However, these methods depend on constrained multi-view images, making hair creation costly and challenging in less controlled settings. Recently, 3D hair generation approaches [92, 13] have emerged using GANs or parametric models, but they struggle with controllable generation. The most relevant work to ours is HAAR [65], which pioneers text-to-strand by training a text-conditioned hair map diffusion model. However, HAAR relies on a large-scale paired dataset including 9825 haircuts and prompts with limited diversity, restricting its ability to generate novel hairstyles beyond the training data. Moreover, HAAR overlooks hair texture and geometry that adapts to specific head shapes, further limiting its application range.
Compared to previous methods, StrandHead generates 3D heads featuring strand-level hairstyles. Instead of requiring large-scale hair-text paired data, we achieve realistic and diverse 3D hair strand generation by using 2D/3D human-centric priors. We summarize the main differences between our methods and related works in Tab. 1.
Figure 2: StrandHead includes three stages: (a) We first create a FLAME-aligned 3D bald head using the improved HumanNorm [17]. (b) Next, we introduce a differentiable prismatization algorithm to enable human-specific geometry-aware 2D diffusion models to supervise hair shape modeling. Additionally, two losses inspired by 3D hair geometric priors are applied to further regularize the hair geometry. (c) Finally, we use a human-specific normal-conditioned 2D diffusion model to generate lifelike hair textures.
@@ 3.1 Preliminaries
FLAME [26] is a 3D parametric head model that represents the human head shape, pose, and expression using a compact set of parameters. Given a set of shape parameters $\beta$, pose parameters ${\theta}_{\text{pose}}$, and expression parameters $\psi_{\text{exp}}$, FLAME can create a 3D head mesh ${M}$.
HumanNorm [17] is a text-driven, high-quality 3D head generation method. In specific, it models the geometry and texture of the human head using DMTet [60] and a texture field. The alignment between the text $y$ and the optimized representation $\theta$ is achieved using the following SDS loss [47] with T2I diffusion models $\phi$:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\phi}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],\vskip-8.5359pt$ (1)
where $x=g(\theta)$ is the image rendered from $\theta$ by a differentiable renderer $g$, $t$ is the time step, $x_{t}=x+\epsilon$ is a noised version of $x$, ${\epsilon}_{\phi}(x_{t};y,t)$ is the denoised image, and $w(t)$ is a weighting function. HumanNorm employs human-specific diffusion models fine-tuned on high-quality human mesh data to replace the general diffusion model, enhancing the fidelity of the generated human heads.
Neural Scalp Textures (NST) [55] is an efficient hair representation where each pixel stores a feature vector conveying the shape information of a single strand at the corresponding scalp location. Using a pre-trained hair strand generator [55] $G$, the low-dimensional 2D neural scalp texture $T$ can be decoded into high-dimensional 3D strand polylines $S=\{s^{i}\}_{i=1}^{N_{s}}$, where $N_{s}$ is the number of strands and $s^{i}=\{p^{i}_{j}\}_{j=1}^{N_{p}}$ consists of $N_{p}$ 3D points. This process is formulated as follows: $S=G(T)$.
@@ 3.2 Overview
Given a text prompt, StrandHead aims to create a realistic 3D head avatar with strand-based hair without relying on large-scale 3D-text paired data. Fig. 2 provides an overview of our three-stage pipeline. First, we generate a FLAME-aligned 3D bald head for accurate hair initialization using human-specific diffusion models (Sec. 3.3). In the subsequent stages, we model the reasonable geometry and realistic texture of 3D hair strands by leveraging rich 2D/3D human-centric priors (Sec. 3.4 and Sec. 3.5).
@@ 3.3 Bald Head Generation
To achieve accurate hair initialization and optimize 3D hair strands in subsequent steps using human-specific 2D generative priors, we first need to obtain a reasonable and semantic-aligned 3D bald head from the text (Fig. 2-(a)). To achieve this goal, we enhance HumanNorm [17] by incorporating an evolving FLAME [26] model while optimizing the bald head DMTet $\theta_{\text{h}}$. This provides accurate semantic information and prevents unnatural geometry. Please refer to the Supp. Mat. for details on bald head generation.
3.4 Hair Geometry Generation
With a FLAME-aligned bald head, we can generate 3D hair in the scalp area. Specifically, we first initialize the haircut based on hairstyle descriptions, and then further sculpt the 3D hair shape utilizing 2D/3D human-centric priors.
Hair Initialization. To achieve reasonable and diverse hair initialization, we first utilize ChatGPT [43] to select the 20 most representative hairstyles from the USC-HairSalon Dataset [15]. We then optimize neural scalp textures to fit the selected hairstyle using the following loss function: